OpenStudy (anonymous):

y + 3 = 2x y + 4 = 3x If you use the comparison method to solve the given system, what is the resulting equation? y + 3 = y + 4 2x - 3 = 3x - 4 2x + 3 = 3x + 4

OpenStudy (anonymous):

solve for y, then substitute the y into one another y+3 =2x --> y=2x-3 y+4 =3x --> y= 3x-4 2x-3=3x-4

OpenStudy (anonymous):

multiply the first term by - and the 2nd term by +

OpenStudy (anonymous):

you'll get -y-3=-2x y+4=3x

OpenStudy (anonymous):

now sub the 1st term from the 2nd

OpenStudy (anonymous):

you'll getx=1 so put x=1 in equation 1 you'll get y+3=2(1) y+3=2 y=2-3 y=-1

OpenStudy (anonymous):

now put x and y value in the 2nd equation -1+4=3(1) -1+4=3 3=3

OpenStudy (anonymous):

we got same answer on both sides so its right